Assume that an industry is perfectly competitive. Each firm must hire a manager, and there exists only 50 managers that display extraordinary talent. There is an unlimited supply of managers with average talent. The long run total cost function of the firms run by exceptional managers is LTCE=200+Q2. The long run total cost function of the firms managed by average managers is LTCA=200+2Q2. If market demand for this good is described by Qd=8000-100p, how much economic rent will each extraordinarily talented manager generate for her firm?
Answer #1

Economic rent is the extra money or payment that we receive

Economic rent of extra ordinary manager = Profit of extra ordinary manager

b) For extra ordinary manager

Qd = 8000-100P

100P = 8000-Q

P = 80-0.01Q

TR = pxQ = Total revenue

TR = (80-0.01Q)Q = 80Q-0.01Q2

Marginal revenue = MR = d(TR)/dQ = d(80-0.01Q2)/dQ

MR = 80-0.02Q

TC = 200+Q2

Marginal cost = MC = d(200+Q2)/dQ = 2Q

MR = MC

80-0.02Q = 2Q

Q = 40 (Approximately)

P = 80-0.01*40= 79.6

TR = 79.6*40 = 3184

TC = 200+40*40 = 1800

Profit = 1384